# TrialBench: Multi-Modal AI-Ready Clinical Trial Datasets

### v2

@@ Abstract
Clinical trials are pivotal for developing new medical treatments but typically carry risks such as patient mortality and enrollment failure that waste immense efforts spanning over a decade. Applying artificial intelligence (AI) to predict key events in clinical trials holds great potential for providing insights to guide trial designs. However, complex data collection and question definition requiring medical expertise have hindered the involvement of AI thus far. This paper tackles these challenges by presenting a comprehensive suite of 23 meticulously curated AI-ready datasets covering multi-modal input features and 8 crucial prediction challenges in clinical trial design, encompassing prediction of trial duration, patient dropout rate, serious adverse event, mortality rate, trial approval outcome, trial failure reason, drug dose finding, design of eligibility criteria. Furthermore, we provide basic validation methods for each task to ensure the datasets’ usability and reliability. We anticipate that the availability of such open-access datasets will catalyze the development of advanced AI approaches for clinical trial design, ultimately advancing clinical trial research and accelerating medical solution development. The curated dataset are available at https://huggingface.co/datasets/ML2Healthcare/ClinicalTrialDataset and the validation frameworks are available at https://github.com/ML2Health/ML2ClinicalTrials/tree/main/AI4Trial.
1 Background & Summary
@@
The clinical trial process is an essential step in developing new treatments (e.g., drugs or medical devices). It involves evaluating their safety, appropriate dosage, and effectiveness in treating specific diseases in the human body. However, these exploratory trials, often spanning three or four phases, have a high failure rate [1, 2]. Compounding the issue, clinical trials are known for being time-consuming, labor-intensive, and costly. Clinical development programs containing the set of phase 1-3 trials typically span 7-11 years, cost an average of 2 billion USD, and achieve approval rates of only around $15\%$ [3]. Clinical trials are inherently risky as they explore “new” treatments, while artificial intelligence (AI) is particularly well-suited for making accurate estimates to reduce risk since AI excels at identifying patterns, including those not previously known to humans.
@@
Years of clinical trials have generated a vast amount of multi-modal data, encompassing aspects such as inclusion/exclusion criteria designs, adverse event statistics, and patient enrollment results. Such extensive data offers a robust foundation for developing advanced artificial intelligence algorithms. However, identifying key clinical trial challenges and effectively leveraging the complex variables within this data require a blend of deep medical knowledge and AI expertise. This complexity has hindered skilled AI experts from fully utilizing the data.
@@
The ClinicalTrials.gov websiteaaahttps://clinicaltrials.gov/ provides comprehensive information on clinical trials, including study protocols, participant eligibility criteria, and study results, making it a valuable resource for AI engineers and medical professionals. This centralized repository covers more than 480,000 clinical trial records (as of Feb 2024) from all 50 US states and international trials from 221 countries. However, identifying key clinical trial challenges suitable for AI solutions and selecting appropriate variables for different challenges remain problematic for data scientists who lack relevant background knowledge.
@@
To facilitate cross-disciplinary research and fully leverage the expertise of data scientists and AI experts [4, 5], this paper identifies 8 key critical clinical trial challenges. It organizes 23 corresponding AI-ready datasets to support their involvement in these tasks. The data, representing clinical trials registered before February 16, 2024, were collected from ClinicalTrials.gov. We extracted elements and attributes from the XML records of each clinical trial and converted them into tabular data formats, which are better suited for processing by AI models, including deep learning models. Additionally, we transformed some features into more informative forms; for example, converting health condition information into ICD-10 codes. We also contextualize our input variables with valuable data from DrugBank [6] and TrialTrovebbbhttps://pharmaintelligence.informa.com/products-and-services/data-and-analysis/trialtrove to depict a comprehensive set of information for clinical trial AI.
@@
The datasets in TrialBench lend themselves to the study of the following open questions in artificial intelligence:
  1. 1.
    Trial Duration Forecasting: The extensive duration of clinical trials, averaging 7 to 11 years, poses a significant challenge in the development of new treatments [12]. Moreover, research indicates that most trials encounter unforeseen delays attributed to strategic challenges, commercial barriers, operational issues, and high toxicity levels [13, 2, 14]. Accurate prediction of trial duration helps pharmaceutical companies better estimate costs, understand how many sites to include to accelerate if necessary, and plan for market launches.
  2. 2.
    Patient Dropout Rate Prediction: Previous research has pointed out that approximately 30% of participants eventually drop out of trials [15]. If patients drop out without completing the primary outcomes, the investment in enrollment will be wasted. More importantly, patient dropout can introduce bias in the study’s conclusions [16]. Thus, predicting patient dropout is crucial for the success of clinical trials, as it helps to minimize follow-up loss and reduce this potential bias. Previous researchers have identified several factors influencing dropout rates, such as age, sex, and education levels [17], providing a solid foundation for using artificial intelligence to predict patient dropout rates during the trial design stage.
  3. 3.
    Serious Adverse Event Prediction: Serious adverse events (SAEs) [18] are critical in assessing the safety profile of new treatments. SAEs can lead to significant trial modifications, including early termination, impacting the development timeline and financial investment [19]. Accurate prediction of SAE can determine patient population size and how long you need to monitor patients, allowing for better risk management and resource allocation.
  4. 4.
    Mortality Event Prediction: When serious adverse events reach a critical level, unsafe treatments or severe diseases may result in fatalities. Unexpected high mortality rates can raise ethical concerns and necessitate comprehensive safety reviews [20]. Predicting mortality events is essential in clinical trial design to ensure patient safety and maintain ethical standards. Mortality events not only can lead to the early termination of trials but also significantly impact the development timeline and financial resources. Developing artificial intelligence algorithms to identify clinical trials at risk of mortality events can significantly enhance risk management and resource allocation.
  5. 5.
    Trial Approval Prediction: Clinical trial suffers from low approval rate [21]. It would save a great amount of funding, time, and other resources if we could forecast the approval rate of the clinical trial before it starts and circumvent the trials with low approval expectations. That is to say, an accurate clinical trial approval model can help clinical decision-makers prioritize trials that are more likely to succeed and promote profit rate.
  6. 6.
    Trial Failure Reason Identification: Failures in clinical trials are commonplace, yet they lead to wasted millions of dollars and years of time, thereby delaying the development of new treatments [22]. Therefore, when designing clinical trials, it is particularly important to anticipate and mitigate risks such as efficiency issues, safety concerns, and poor enrollment. By understanding the root risks of failures in advance, trialists can make informed decisions to enhance design.
  7. 7.
    Eligibility Criteria Design: The design of inclusion and exclusion eligibility criteria is one of the most crucial aspects of clinical trial design, significantly impacting the success or failure of the trial. Both overly restrictive and excessively lenient criteria [23] can lead to low enrollment rates in clinical trials [24, 25] or result in insufficient conclusions. Currently, the design of these criteria primarily relies on the experience of physicians, which can sometimes lack objectivity. Utilizing existing natural language processing technologies, particularly large language models (LLMs), to automate the design of inclusion and exclusion eligibility criteria presents a promising advantage.
  8. 8.
    Drug Dose Finding: In Phase II trials, controlling drug dosage is crucial for ensuring safety while maintaining medication efficiency [26]. However, drug dosage control largely relies on the experience of physicians, which can be subjective. Artificial intelligence algorithms offer a promising approach to predict and optimize drug dosages based on drug SMILES and target diseases. By accurately predicting the appropriate drug dosage, clinical trials can enhance patient safety, improve treatment efficacy, and potentially reduce the duration and costs associated with trial phases.
@@
Fig. 1 illustrates the TrialBench platform, containing 8 well-defined clinical trial design tasks. The TrialBench platform provides 23 corresponding AI-ready datasets, implemented evaluation metrics, and baseline models. AI experts can easily access the datasets and targets to develop advanced models, evaluate models on specific metrics, and compare them against baseline models for reference.
2 Methods
2.1 AI-solvable Clinical Trial Task Definitions
@@
In this paper, we identify nine AI-solvable clinical trial tasks. For each task, we elaborate on its background, explain how it would help clinical trial design and management, curate the dataset, evaluate the performance of well-known artificial intelligence methods, and report the empirical results. Table 1 summarizes and compares all the AI-solvable clinical trial tasks and corresponding datasets. We provide the following three aspects for each learning task: (1) Background. Background of the learning task. (2) Definition. A formal definition of the learning task (input feature and output). (3) Broad impact. The broader impact of advancing real clinical trials on the task.
@@
Problem AI Task Input Modality Input Output # Data
trial duration forecasting Regression all 5 modalities See Fig. 6 trial duration (e.g., 3.2 years) 141,940
patient dropout event forecasting Classification/ Regression all 5 modalities See Fig. 7 patient dropout event (0/1) / rate [0,1) 62,058
serious adverse event forecasting Classification all 5 modalities See Fig. 7 adverse event (0/1) 31,306
mortality event prediction Classification all 5 modalities See Fig. 7 mortality rate (0/1) 31,306
trial approval forecasting Classification all 5 modalities See Fig. 8 trial approval label (0/1) 43,202
trial failure reason identification Classification all 5 modalities See Fig. 9 trial failure reasons (4 categories) 41,369
eligibility criteria design Generation MeSH, SMILES, ICD-10, Texts See Fig. 10 eligibility criteria (natural language) 136,443
drug dose finding Classification SMILES, MeSH SMILES & intervention MeSH drug dosage (4 categories) 12,790
2.1.1 Trial Duration Prediction
@@
Definition. This task focuses on predicting trial duration (time span from the enrollment of the first participant to the conclusion of the study) based on multi-modal trial features such as eligibility criteria, target disease, etc.
@@
2.1.2 Patient Dropout Forecasting
Background. Clinical trials often suffer from high patient dropout rates, which can compromise the validity of the results and lead to increased costs and delays.
Definition. This task aims to predict the patient dropout rate (percentage) of the clinical trial based on multi-modal trial features such as eligibility criteria, target disease, and so on.
Broad impact. Predicting patient dropout rates in clinical trials holds significant promise for improving the efficiency and effectiveness of drug development processes. Predicting patient dropout rates can significantly enhance the efficiency of clinical trials. High dropout rates often necessitate the recruitment of additional participants to meet the required sample size, which can be both time-consuming and costly.
2.1.3 Serious Adverse Rate Prediction
@@
Definition. The task targets forecasting the probability of serious adverse effects given multi-modal clinical trial features such as drug molecule, target disease, eligibility criteria, etc.
@@
2.1.4 Mortality Rate Prediction
@@
Background. The mortality rate in a clinical trial refers to the proportion of participants who die during the study. The mortality rate is an important measure used to assess the safety and potential risks associated with a treatment or intervention being tested in the trial.
@@
Definition. The task targets forecasting the probability of mortality rate given multi-modal clinical trial features such as drug molecule, target disease, eligibility criteria, etc.
@@
Broad impact. Accurately predicting the mortality rate of a clinical trial significantly enhances patient safety by identifying potential risks early, allowing for timely interventions. This leads to more efficient trial designs, optimizing resource allocation and reducing costs. Furthermore, it accelerates the drug development process, bringing effective treatments to market faster, and increases compliance with regulatory standards, thereby building public trust and ethical standards in clinical research.
2.1.5 Trial Approval Prediction
@@
Background. Clinical trial approval refers to whether a drug can pass a certain phase of clinical trial, which is the most important outcome of a clinical trial. It is a binary variable.
@@
Definition. This task aims to predict the probability of trial approval given multi-modal trial features such as drug molecule, disease code, and eligibility criteria.
@@
Broad impact. Predicting trial approval can significantly enhance the efficiency and success rates of drug development. By accurately forecasting which drugs are likely to pass clinical trial phases, companies can focus their resources on the most promising candidates, reducing wasted time and money on less viable options. This targeted approach can accelerate the development of effective treatments, bringing them to market faster and improving patient outcomes. Additionally, reliable approval predictions can streamline regulatory processes and increase investor confidence in the pharmaceutical industry.
2.1.6 Trial Failure Reason Identification
@@
Background. Clinical trials usually fail due to a couple of reasons [30]: (1) business decision (e.g., lack of funding, company strategy shift, pipeline reorganization, drug strategy shift); it is challenging to predict business decision, so we do not involve these trial in our dataset; (2) Poor enrollment. Insufficient enrollment can compromise the statistical power of the study, making it difficult to detect a significant effect of the drug. Also, poor enrollment can lead to delays in the trial timeline and increased costs, as more resources are required to recruit additional participants. (3) Safety. Unexpected adverse reactions or side effects can occur, posing significant risks to participants’ health. This can lead to the trial being halted or terminated. (4) Efficacy (effectiveness). In the trial, we expect the tested drug to outperform the standard treatment in curing the target disease. Thus, efficacy (effectiveness) is typically required.
@@
Definition. Given clinical trial features, the goal of this task is to leverage the AI model to classify it into one of these four categories, including (1) successful trials, (2) failure due to poor enrollment, (3) failure due to drug safety issue; (4) fail due to lack of efficacy. It is a multi-category classification problem.
@@
2.1.7 Eligibility Criteria Design
Background. To achieve statistically significant results, a clinical trial must meet its target sample size [31]. Insufficient patient numbers can lead to underpowered studies, which may fail to demonstrate the effectiveness of a treatment or may miss important safety information. Eligibility criteria are essential to patient recruitment [32]. They describe the patient recruitment requirements in unstructured natural language. Eligibility criteria comprise multiple inclusion and exclusion criteria, which specify what is desired and undesired when recruiting patients. Each individual criterion is usually a natural language sentence.
@@ 2.1.7 Eligibility Criteria Design
2.1.8 Drug Dose Finding
Background. One of the primary goals of clinical trials is to determine the drug dose. Determining the correct dosage of a drug is crucial to ensure its effectiveness in treating a particular condition. In the early stages of drug development, predicting the optimal dosage is essential for designing clinical trials [26, 33].
Definition. This task aims to predict drug dosage based on drug molecular structure and target disease.
@@ 2.1.8 Drug Dose Finding
2.2 Data Acquisition
@@
  • ClinicalTrials.gov. ClinicalTrials.gov is a publicly accessible database maintained by the U.S. National Library of Medicine (NLM) at the National Institutes of Health (NIH). It provides detailed information about clinical trials conducted around the world, including those funded by public and private entities. Each clinical trial in ClinicalTrials.gov is provided as an XML file, which we parse to extract relevant variables. For each trial, we retrieve the NCT ID (unique identifiers for each clinical study), disease names, associated drugs, title, summary, trial phase, eligibility criteria, results of statistical analyses, and other details. Some of these features are not always available. For example, observational clinical trials do not involve treatment and drugs.
  • DrugBank. DrugBank [6] (https://www.drugbank.com/) is a comprehensive, freely accessible online database that provides detailed information about drugs and their biological targets. It integrates chemical, pharmacological, and pharmaceutical data with comprehensive drug target information, making it a valuable resource for researchers, healthcare professionals, and students in the fields of drug discovery, pharmacology, and medicinal chemistry.
  • TrialTrove. TrialTroveccchttps://pharmaintelligence.informa.com/products-and-services/data-and-analysis/trialtrove is a comprehensive database and intelligence platform designed to provide detailed information and analysis on clinical trials across the pharmaceutical and biotechnology industries. TrialTrove serves as a critical resource for professionals involved in clinical development, competitive intelligence, and market analysis. We obtain the trial outcomes of some trials from the released/public subset of the TrialTrove database [34, 35].
  • ICD-10. ICD-10-CM (International Classification of Diseases, 10th Revision, Clinical Modification) is a medical coding system for classifying diagnoses and reasons for visits in U.S. healthcare settings. Diseases are extracted from https://clinicaltrials.gov/ and linked to ICD-10 codes and disease description using Clinical Table Search Service APIdddclinicaltables.nlm.nih.gov and then to CCS codes via hcup-us.ahrq.gov/toolssoftware/ccs10/ccs10.jsp.
@@
2.3 Dataset Curation and Feature Organization
We apply a series of selection filters to ensure the selected trials have high-quality. There are hundreds of multi-modal features in ClinicalTrials.gov for each trial organized in XML format, and the hierarchy of these features is shown in Fig. 5. We only leverage the features that are available before trials start and remove the remaining features. Different tasks rely on different subsets of features. Based on clinical trial knowledge, we manually select the appropriate features for various tasks. In addition, we also remove features whose values are identical or all null across different trials. Following are the additional selection criteria for each task.
@@ 2.4 Data Annotation
  • Trial failure reason identification: For some of the terminated trials, ClinicalTrials.gov provides a “why stopped” tag that uses natural language to describe the failure reason. We use OpenAI ChatGPT API eeehttps://openai.com/index/openai-api/ to automatically convert into three categories of failure reason, including (1) poor enrollment; (2) drug safety issue; (3) lack of efficacy (in treating the target disease). In using ChatGPT, the prompt and instruction are shown below, and we required ChatGPT to complete the “reasons” part:
@@
  • Drug dose finding: One aim of phase-II clinical trials is to determine the dosage of the drug. ClinicalTrials.gov presents the drug dosage information of some trials in natural language. We use OpenAI ChatGPT APIfootnotemark: to extract the label from natural language, the prompt is shown below.
@@ 2.5 Data Partitioning
3 Data Records
3.1 Data Overview
Our primary data source is the clinicalTrials.gov website (https://clinicaltrials.gov/), which serves as a publicly accessible resource for clinical trial information. Supported by the U.S. National Library of Medicine, this database encompasses over 420,000 clinical trial records, spanning all 50 U.S. states and 221 countries worldwide. The number of recorded trials would grow rapidly with time, as shown in Figure 2 (a). Table 2 reports some essential statistics of the curated datasets, including the number of involved trials, drugs, diseases, and proportion of interventional trials. There are hundreds of multi-modal features in ClinicalTrials.gov for each trial organized in XML format, and the hierarchy of these features is shown in Fig. 5. Table 5 demonstrates a real clinical trial example. Originally organized in an XML hierarchy format, we selected relevant features based on the challenges of each task and re-organized them into a tabular data format. Notably, in addition to categorical and numerical tabular features, some of these features may include free text, graph data, and other complex types.
@@ 3.1 Data Overview
Here, we review some essential features for better utilization of our datasets. Notably, some trials have missing features; for example, certain incomplete trials lack a completed date and outcome.

### v3

# TrialBench: Multi-Modal AI-Ready Datasets for Clinical Trial Prediction Datasets available at: https://huyjj.github.io/Trialbench/, accepted by Nature Scientific Data

@@ Abstract
Clinical trials are pivotal for developing new medical treatments but typically carry risks such as patient mortality and enrollment failure that waste immense efforts spanning over a decade. Applying artificial intelligence (AI) to predict key events in clinical trials holds great potential for providing insights to guide trial designs. However, complex data collection and question definition requiring medical expertise have hindered the involvement of AI thus far. This paper tackles these challenges by presenting a comprehensive suite of 23 meticulously curated AI-ready datasets covering multi-modal input features and 8 crucial prediction challenges in clinical trial design, encompassing prediction of trial duration, patient dropout rate, serious adverse event, mortality rate, trial approval outcome, trial failure reason, drug dose finding, design of eligibility criteria. Furthermore, we provide basic validation methods for each task to ensure the datasets’ usability and reliability. We anticipate that the availability of such open-access datasets will catalyze the development of advanced AI approaches for clinical trial design, ultimately advancing clinical trial research and accelerating medical solution development.
Background & Summary
@@
The clinical trial process is an essential step in developing new treatments (e.g., drugs, vaccines, or medical devices), serving as the bridge between scientific discovery and real-world medical application. Clinical trials are designed to systematically evaluate the safety, efficacy (in treating specific diseases), dosage, and overall impact of these treatments on human bodies [1]. The basic steps of a clinical trial typically include: (1) planning and design, where researchers define the study objectives, eligibility criteria, and determine the treatment protocol; (2) recruitment and screening, where eligible participants are enrolled and baseline health data is collected; (3) intervention and monitoring, where participants receive treatment or placebo, and their health outcomes are closely monitored; (4) data analysis, where results are analyzed to determine the treatment’s safety and efficacy; (5) conclusion, findings are reported and, if successful, submitted for regulatory approval. Typically conducted in multiple phases (Phase 1 to Phase 4, approved by FDA after passing Phase 3), these trials begin with small-scale studies to assess safety and dosage (Phase 1, 20-80 healthy volunteers), expand to evaluate efficacy and side effects in larger populations (Phase 2 and 3, 100-300 and 300-3,000 patients respectively), and continue into post-marketing surveillance to monitor long-term outcomes (Phase 4, several thousand to tens of thousands of patients) [2]. However, these exploratory trials have a high failure rate [3, 4]. Compounding the issue, clinical trials are known for being time-consuming, labor-intensive, and costly. Clinical development programs containing the set of phase 1-3 trials typically span 7-11 years, cost an average of 2 billion USD, and achieve approval rates of only around $15\%$ [5]. Clinical trials are inherently risky as they explore “new” treatments, while artificial intelligence (AI) is particularly well-suited for making accurate estimates to reduce risk since AI excels at identifying patterns, including those not previously known to humans [6] .
@@
Years of clinical trials have generated a vast amount of multi-modal data [7] , encompassing aspects such as inclusion/exclusion criteria designs, adverse event statistics, and patient enrollment results. Such extensive data offers a robust foundation for developing advanced AI algorithms [8] . However, identifying key clinical trial challenges and effectively leveraging the complex variables within this data require a blend of deep medical knowledge and AI expertise. This complexity has hindered skilled AI experts from fully utilizing the data.
@@
The ClinicalTrials.gov website (https://clinicaltrials.gov/) provides comprehensive information on clinical trials, including study protocols, participant eligibility criteria, and study results, making it a valuable resource for AI engineers and medical professionals. This centralized repository covers more than 480,000 clinical trial records (as of Feb 2024) from all 50 US states and international trials from 221 countries. However, identifying key clinical trial challenges suitable for AI solutions and selecting appropriate variables for different challenges remain problematic for data scientists who lack relevant background knowledge.
@@
To facilitate cross-disciplinary research and fully leverage the expertise of data scientists and AI experts [9, 10], this paper identifies 8 key critical clinical trial challenges. It organizes 23 corresponding AI-ready datasets to support their involvement in these tasks. The data, representing clinical trials registered before February 16, 2024, were collected from ClinicalTrials.gov. We extracted elements and attributes from the XML records of each clinical trial and converted them into tabular data formats, which are better suited for processing by AI models, including deep learning models. Additionally, we transformed some features into more informative forms; for example, converting health condition information into ICD-10 codes. We also enrich our data with valuable information from DrugBank (e.g., drug molecular structures and pharmaceutical properties as feature) [11] and TrialTrove (e.g., trial approval information as groundtruth) (https://pharmaintelligence.informa.com/products-and-services/data-and-analysis/trialtrove) to depict a comprehensive set of information for clinical trial AI. Our organized datasets are available at: https://huyjj.github.io/Trialbench/.
@@
Fig. 1 illustrates the TrialBench platform, containing 8 well-defined clinical trial design tasks. The TrialBench platform provides 23 corresponding AI-ready datasets across these 8 tasks, implemented evaluation metrics, and baseline models. AI experts can easily access the datasets and targets to develop advanced models, evaluate models on specific metrics, and compare them against baseline models for reference.
Methods
AI-solvable Clinical Trial Task Definitions
@@ Methods
In this paper, we identify 8 AI-solvable clinical trial tasks. For each task, we elaborate on its background, explain how it would help clinical trial design and management, curate the dataset, evaluate the performance of well-known artificial intelligence methods, and report the empirical results. Table 1 summarizes and compares all the AI-solvable clinical trial tasks and corresponding datasets. We provide the following three aspects for each learning task: (1) Background. Background of the learning task. (2) Definition. A formal definition of the learning task (input feature and output). (3) Broad impact. The broader impact of advancing real clinical trials on the task.
@@
Problem AI Task Input Modality Input & # features Output # Data
trial duration forecasting Regression all 5 modalities See Fig. S2 & 35 trial duration (e.g., 3.2 years) 141,940
patient dropout event forecasting Classification/ Regression all 5 modalities See Fig. S3 & 38 patient dropout event (0/1) / rate [0,1) 62,058
serious adverse event forecasting Classification all 5 modalities See Fig. S3 & 38 adverse event (0/1) 31,306
mortality event prediction Classification all 5 modalities See Fig. S3 & 38 mortality rate (0/1) 31,306
trial approval forecasting Classification all 5 modalities See Fig. S4 & 35 trial approval label (0/1) 43,202
trial failure reason identification Classification all 5 modalities See Fig. S5 & 32 trial failure reasons (4 categories) 41,369
eligibility criteria design Generation MeSH, SMILES, ICD-10, Texts See Fig. S6 & 25 eligibility criteria (natural language) 136,443
drug dose finding Classification SMILES, MeSH SMILES & intervention MeSH & 2 drug dosage (4 categories) 12,790
Trial Duration Prediction
@@
Definition. This task focuses on predicting trial duration (time span from the enrollment of the first participant to the conclusion of the study) based on multi-modal trial features such as eligibility criteria, target disease, etc. It is formulated as a regression task.
@@
Patient Dropout Forecasting
Background. Previous research has pointed out that approximately 30% of participants eventually drop out of trials [18], which can compromise the validity of the results and lead to increased costs and delays.
Definition. This task aims to predict the patient dropout rate (percentage) of the clinical trial based on multi-modal trial features such as eligibility criteria, target disease, etc. It is formulated as a dual-objective task: a binary classification task for predicting dropout occurrence and a regression task for forecasting the dropout rate.
Broad impact. Predicting patient dropout rates in clinical trials holds significant promise for improving the efficiency and effectiveness of drug development processes. Predicting patient dropout rates can improve the efficiency of clinical trials. High dropout rates often necessitate the recruitment of additional participants to meet the required sample size, which can be both time-consuming and costly.
Serious Adverse Rate Prediction
@@
Definition. The task targets forecasting the probability of serious adverse effects given multi-modal clinical trial features such as drug molecule, target disease, eligibility criteria, etc. It is formulated as a binary classification problem.
@@
Mortality Rate Prediction
@@
Background. The mortality rate in a clinical trial refers to the proportion of participants who die during the study. When serious adverse events reach a critical level, unsafe treatments or severe diseases may result in fatalities. Unexpected high mortality rates can raise ethical concerns and necessitate comprehensive safety reviews [21]. The mortality rate is an important measure used to assess the safety and potential risks associated with a treatment or intervention being tested in the trial.
@@
Definition. The task targets forecasting the probability of mortality rate given multi-modal clinical trial features such as drug molecule, target disease, eligibility criteria, etc, which is formulated as a binary classification problem.
@@
Broad impact. Accurately predicting the mortality rate of a clinical trial enhances patient safety by identifying potential risks early, allowing for timely interventions. This leads to more efficient trial designs, optimizing resource allocation and reducing costs. Furthermore, it accelerates the drug development process, bringing effective treatments to market faster, and increases compliance with regulatory standards, thereby building public trust and ethical standards in clinical research.
Trial Approval Prediction
@@
Background. Clinical trial approval refers to whether a drug can pass a certain phase of clinical trial, which is the most important outcome of a clinical trial. Recent investigations suggest that clinical trial suffers from low approval rate [22].
@@
Definition. This task aims to predict the probability of trial approval given multi-modal trial features such as drug molecule, disease code, and eligibility criteria. It is formulated as a binary classification problem.
@@
Broad impact. Predicting trial approval can enhance the efficiency and success rates of drug development. By accurately forecasting which drugs are likely to pass clinical trial phases, companies can focus their resources on the most promising candidates, reducing wasted time and money on less viable options. This targeted approach can accelerate the development of effective treatments, bringing them to market faster and improving patient outcomes. Additionally, reliable approval predictions can streamline regulatory processes and increase investor confidence in the pharmaceutical industry.
Trial Failure Reason Identification
@@
Background. Clinical trials usually fail due to a couple of reasons [23]: (1) business decision (e.g., lack of funding, company strategy shift, pipeline reorganization, drug strategy shift); it is challenging to predict business decision, so we do not involve these trials in our dataset; (2) Poor enrollment. Insufficient enrollment can compromise the statistical power of the study, making it difficult to detect a significant effect of the drug. Also, poor enrollment can lead to delays in the trial timeline and increased costs, as more resources are required to recruit additional participants. (3) Safety. Unexpected adverse reactions or side effects can occur, posing significant risks to participants’ health. This can lead to the trial being halted or terminated. (4) Efficacy. In the trial, we expect the tested drug to outperform the standard treatment in curing the target disease. Thus, efficacy is typically required.
@@
Definition. Given clinical trial features, the goal of this task is to leverage the AI model to classify it into one of these four categories, including (1) successful trials, (2) failure due to poor enrollment, (3) failure due to drug safety issue; (4) fail due to lack of efficacy. It is a multi-category (4 categories) classification problem.
@@
Eligibility Criteria Design
Background. To achieve statistically significant results, a clinical trial must meet its target sample size [24]. Insufficient patient numbers can lead to underpowered studies, which may fail to demonstrate the efficacy of a treatment or may miss important safety information. Eligibility criteria are essential to patient recruitment [25]. They describe the patient recruitment requirements in unstructured natural language. Eligibility criteria comprise multiple inclusion and exclusion criteria, which specify what is desired and undesired when recruiting patients. Each individual criterion is usually a natural language sentence.
@@
Drug Dose Finding
Background. One of the primary goals of clinical trials is to determine the drug dose. Determining the correct dosage of a drug is crucial to ensure its efficacy in treating a particular condition. In the early stages of drug development, predicting the optimal dosage is essential for designing clinical trials [26, 27].
Definition. This task aims to predict drug dosage based on drug molecular structure and target disease, which is formulated as an ordinal classification problem.
@@
Raw Data: ClinicalTrials.gov
Our primary data source is the clinicalTrials.gov website (https://clinicaltrials.gov/), which serves as a publicly accessible resource for clinical trial information. Supported by the U.S. National Library of Medicine, this database encompasses over 420,000 clinical trial records, spanning all 50 U.S. states and 221 countries worldwide. The number of recorded trials would grow rapidly with time, as shown in Figure 2 (a). Table 3 reports some essential statistics of the curated datasets, including the number of involved trials, drugs, diseases, and proportion of interventional trials. There are hundreds of multi-modal features in ClinicalTrials.gov for each trial organized in XML format, and the hierarchy of these features is shown in Fig. S1. Table 2 demonstrates a real clinical trial example.
Table 2: A real example of a clinical trial record.
Feature Descriptions
NCTID NCT00610792
disease Ovarian Cancer
phase II
study type interventional
drug bortezomib and pegylated liposomal doxorubicin
start date July 2006
completed date September 2009
sponsor Millennium Pharmaceuticals, Inc.
outcome withdrawn
Data Acquisition
@@
  • ClinicalTrials.gov. ClinicalTrials.gov is a publicly accessible database maintained by the U.S. National Library of Medicine (NLM) at the National Institutes of Health (NIH). It provides detailed information about clinical trials conducted around the world, including those funded by public and private entities. Each clinical trial in ClinicalTrials.gov is provided as an XML file, which we parse to extract relevant variables. For each trial, we retrieve the NCT ID (unique identifiers for each clinical study), disease names, associated drugs, title, summary, trial phase, eligibility criteria, results of statistical analyses, other details, and then integrate into our data. Some of these features are not always available. For example, observational clinical trials do not involve treatment and drugs.
  • DrugBank. DrugBank [11] (https://www.drugbank.com/) is a comprehensive, freely accessible online database that provides detailed information about drugs and their biological targets. We extract the drug molecular structures and pharmaceutical properties from DrugBank, which are essential to drug’s safety in human bodies and efficacy in treating certain diseases.
  • TrialTrove. TrialTrove (https://pharmaintelligence.informa.com/products-and-services/data-and-analysis/trialtrove) is a comprehensive database and intelligence platform designed to provide detailed information and analysis on clinical trials across the pharmaceutical and biotechnology industries. TrialTrove serves as a critical resource for professionals involved in clinical development, competitive intelligence, and market analysis. We obtain the trial outcomes of some trials from the released/public subset of the TrialTrove database [28, 29].
  • ICD-10. ICD-10-CM (International Classification of Diseases, 10th Revision, Clinical Modification) is a medical coding system for classifying diagnoses and reasons for visits in U.S. healthcare settings. Diseases are extracted from https://clinicaltrials.gov/ and linked to ICD-10 codes and disease description using Clinical Table Search Service API(clinicaltables.nlm.nih.gov) and then to CCS codes via hcup-us.ahrq.gov/toolssoftware/ccs10/ccs10.jsp.
@@
Dataset Curation and Feature Organization
We apply a series of selection filters to ensure the selected trials have high-quality. There are hundreds of multi-modal features in ClinicalTrials.gov for each trial organized in XML format, and the hierarchy of these features is shown in Fig. S1. We only leverage the features that are available before trials start and remove the remaining features. Different tasks rely on different subsets of features. Based on clinical trial knowledge, we manually select the appropriate features for various tasks. In addition, we also remove features whose values are identical or all null across different trials. Following are the additional selection criteria for each task.
@@
  • Trial failure reason identification: For some of the terminated trials, ClinicalTrials.gov provides a “why stopped” tag that uses natural language to describe the failure reason. We use OpenAI ChatGPT API (https://openai.com/index/openai-api/) to automatically convert into four categories of failure reason, including (1) poor enrollment; (2) drug safety issue; (3) lack of efficacy (in treating the target disease); (4) others (e.g., lack of funding, strategic decision by sponsor). Since the last failure reason ((4) others) is usually not predictable, we perform 4-category classification ((1) success; (2) poor enrollment; (3) drug safety issue (4) lack of efficacy). In using ChatGPT, the prompt and instruction are shown below, and we required ChatGPT to complete the “reasons” part:
@@
  • Drug dose finding: One aim of phase-II clinical trials is to determine the dosage of the drug. ClinicalTrials.gov presents the drug dosage information of some trials in natural language. We use OpenAI ChatGPT API (https://openai.com/index/openai-api/) to extract the label from natural language, the prompt is shown below.
@@
Data Records
Data Overview
Clinical trial records are originally organized in an XML hierarchy format, we selected relevant features based on the challenges of each task and re-organized them into a tabular data format. Notably, in addition to categorical and numerical tabular features, some of these features may include free text, graph data, and other complex types.
Here, we review some essential features in our datasets. Notably, some trials have missing features; for example, certain incomplete trials lack a completed date and outcome.